OpenStudy (anonymous):

A bag of pins has 4 black, 2 brown and 6 blue pins in it. If 3 pins are selected without replacement then determine the probability of getting red, green and blue in that order?

OpenStudy (anonymous):

fff

OpenStudy (anonymous):

probility of red pins = 4/12 probility of green pins = 0 (there are no green pins) probility of blue pins = 6/12

OpenStudy (anonymous):

sorry brown